Circular No. 6223

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) Z. Sekanina, Jet Propulsion Laboratory, writes regarding the items on

IAUC 6216 and 6220: “The fact that the jet did not move much in 24 hr, combined with the meager effect of radiation pressure, implies that particle velocities must be very small (≪ 100 m/s). Since the curvature of the jet cannot be due to radiation pressure, it must be due to the spatial distribution of particle velocity vectors—i.e., the particles must have been ejected from an area far from the pole. Furthermore, the jet’s foreshortening orientation suggests that the projected direction of the spin axis was probably along the line near p.a. 100^{o}–280^{o}; the earth could not be too close to the equatorial plane, or we should see the spiral projected nearly edge-on. If one assumes that the ‘tip’ of the jet near p.a. 290^{o} was the beginning of the emission episode and its ‘top’ (slightly east of north) was its end, occurring at local sunset, one gets a compact set of solutions for the spin-axis position (α ∼ 45^{o}, δ ∼ +10^{o}, obliquity 140^{o}–150^{o}). In any event, the emission began about when the source transited the subsolar meridian. The rotation period is entirely indeterminate, except that it could not be greater than ∼ 10 days, but the emission episode must have lasted very nearly \frac{1}{4} of the rotation period. Assuming a particle velocity of 10 m/s (unforeshortened) on the near side of the jet (relative to the nucleus), the age of the jet (on Aug. 31) might be ∼ 35 days. The available images suggest that this emission episode was not the only one, a conclusion that is consistent with the spiral reported in late July (IAUC 6194). One can infer that C/1995 O1 might be (relatively) CO-rich and, at the same time, dust-rich. If such CO supplies would last until perihelion, then the comet could indeed become very bright. Yet, it does not have to have an excessively large nucleus.”

GRO J1735–27 A. J. Castro-Tirado, Laboratorio de Astrofisica Espacial y Fisica Fun-

damental, Madrid, reports: “We examined the error box of the x-ray transient source GRO J1735–27 (IAUC 6207, 6222) on a Schmidt plate taken by K. Birkle at the German-Spanish Calar Alto Observatory on Aug. 25; nothing obvious down to B ∼ 17 is seen. We also expanded our search to the error box of the Ginga source GS 1734–275 (IAUC 4571, 6212), slightly outside of the BATSE error box; that result was negative as well. Further soft-x-ray observations are encouraged to confirm whether GRO J1735–27 is related to the x-ray transient observed by ROSAT (RX J1735.9–2726, IAUC 6212) and Mir-Kvant/TTM (KS 1732–273, IAUC 6216).”

1995 September 8 Daniel W. E. Green